OpenStudy (anonymous):

What is the approx. volume of the sphere if the surface area is 482.8 mm^2?

OpenStudy (anonymous):

To get the volume, you'll need the radius, right? So, to get the radius from the formula: 4*pi*radius^2 = 482.8, how would you isolate the radius variable?

OpenStudy (anonymous):

Divide 482.8 by 4 and pi?

OpenStudy (anonymous):

Yup, divide by 4pi, let me know what you get.

OpenStudy (anonymous):

38.42?

OpenStudy (anonymous):

Affirmative, now since that is the radius squared, how would we get just the radius?

OpenStudy (anonymous):

Nice vocabulary haha, and I would square root it?

OpenStudy (anonymous):

Yes you would.

OpenStudy (anonymous):

since the volume formula uses radius^3, you can also raise it to the (3/2) power, which means the square root to the power of three.

OpenStudy (anonymous):

6.198?

OpenStudy (anonymous):

Correct, now sub what you know into the volume formula: V = (4/3) * pi * radius^3

OpenStudy (anonymous):

998?

OpenStudy (anonymous):

Looks good to me.

OpenStudy (anonymous):

Also, what is the approx. surface area of the sphere?

OpenStudy (anonymous):

Well, the circumference = pi * 2 * radius, so solve for the radius and you should be set. LMK if you run into a roadblock.

OpenStudy (anonymous):

LMK?

OpenStudy (anonymous):

let me know

OpenStudy (anonymous):

Wow I am an idiot

OpenStudy (anonymous):

R = 5.9969?

OpenStudy (anonymous):

Yup, kilometers, big sphere. So plug that into: 4 * pi * radius^2.

OpenStudy (anonymous):

451.9

OpenStudy (anonymous):

That's the answer.

OpenStudy (anonymous):

Sweet! Thanks, I don't know what I would do without your help.